OpenStudy (anonymous):

Yuri solved the equation below incorrectly. Explain the step where he went wrong. (see reply bar for equation)

OpenStudy (anonymous):

15b - 1 = 30 Step 1. 15b - 1 - 1 = 30 - 1 Step 2. 15b = 29 Step 3. \[\frac{ 15b }{ 15 } = \frac{ 29 }{ 15 }\] Step 4. \[B = \frac{ 29 }{ 15 }\]

OpenStudy (austinl):

First off, it should be +1 in step 1. There you go.

OpenStudy (austinl):

\(\large{b=\dfrac{31}{15}}\)